This question is from the textbook called integrated algebra chapter 2 section 2.4 and it is number 35.
My Question:
An angle whose supplement is 20 degrees more than twice its compliment.
- I need to find the measure of the given angles.
If you can help me that would be great. Thank you very much.
ISBN-13 :978-0-07-330917-0
ISBN 10: 0-07-330917-6

Found 2 solutions by edjones, ionesco:
Answer by edjones(7311)

angle=x Compliment=90-x Suppliment=180-x
180-x=2(90-x)+20
180-x=180-2x+20
add 2x to both sides. 180+x=200
subtract 180 from both sides x=20 degrees (Ans.)
compliment=70 Suppliment=160
Twice the compliment + 20 = the suppliment
EdJones

Answer by ionesco(12)

I hope I understood:
we have 3 equations:
1st for suplementary angles: x+y=180
2nd for complementary angles: x+z=90 then z=90-x
then
y= 20 + 2z= 20+ 2(90-x)=200-2x then 2x+y=200
so, x+y=180 and 2x+y=200 where x=20 and y=160